OpenStudy (anonymous):

x^3+8x^2=25x+200 I know that the answer is -5,5,-8 but I dont know how to come up with it.

OpenStudy (anonymous):

if i answer correct can you give me a medal?

OpenStudy (anonymous):

\[x^{3}+8x ^{2}-25x-200=0\] \[(x^{3}+8x)+(-25x-200)=0\] factor: \[x^{2}(x+8)-25(x+8)=0\] \[(x^{2}-25)(x+8)(x+8)=0\] \[(x+5)(x-5)(x+8)(x+8)=0\] (x+8) cancel now :\[(x+5)(x-5)(x+8)=0\] x= 5 x= -5 x=-8

OpenStudy (anonymous):

I let you look ,I be back